Please help me with the linear equation y=-4/3x+1
by finding the y coordinates of the points (-6) (-3) and (6)

Answer by rfer(16322)

y=-4/3x+1
y=-4/3*-6+1
y=8+1
y=9
---------------
y=-4/3*-3+1
y=4+1
y=5
-------------
you do the last one.